Hi can you please help me figure this out. I am not sure if I am doing this right.
I am trying to decide if the given number is a solution to the given equation.
5x^2 405; -9
and if
3x + 7 = 21; 4
I am thinking yes to both but I how do I know for sure.
Thanks for all your help :o)

Answer by mananth(16949)

5x^2=408
x^2=81
x= +/- 9
...
3x+7=21
3x=21-7
3x=14
/3
x=14/3
4 not a solution